OpenStudy (anonymous):

derivative of u^8e(-u)

OpenStudy (anonymous):

This one is tricky, and the first step is probably the hardest part. First, recognize that 8*e is a constant. From there, can you think of a way you can rewrite the expression as an exponential with "e" as the base?

OpenStudy (anonymous):

I am sorry I missed typed the equation. the 8 is the exponent of u and the -u is the exponent of e

OpenStudy (anonymous):

Oh, so is the question supposed to be like this? \[u ^{8}e ^{-u}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, whew, that's much easier. The key here is to remember the chain rule. You basically have two expressions that are multiplied together, so apply the chain rule like this: \[\frac{ d }{ dx } (f(x)*g(x)) = f'(x)*g(x) + g'(x)*f(x)\]Do you understand how to apply that formula to this problem?

OpenStudy (anonymous):

I think so. There are examples I can use to help

OpenStudy (anonymous):

Knowing the rule helps alot.

OpenStudy (anonymous):

Ok, I can carry it out a little bit further by plugging in your example into the formula, if that would help.

OpenStudy (anonymous):

That would be aswesome

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Ok, so with your example,\[f(x) = u ^{8}\] and \[g(x) = e^{-u}\]and since they are multiplied together, \[\frac{ d }{ dx }(f(x)*g(x)) = \frac{d}{dx}(u ^{8})*e ^{-u} + \frac{d}{dx}(e^{-u})*u^8\] Do you think you can take those derivatives?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Sorry, I just realized that I wrote that in a very poor way. I'm so used to writing derivatives with respect to x that I forgot that yours was with respect to u.

OpenStudy (anonymous):

Sorry I had an answer it did not take

OpenStudy (anonymous):

With your numbers plugged in, it should be \[\frac{d}{du}(u^8)*e^{-u} + \frac{d}{du}(e^{-u})*u^8\]

OpenStudy (anonymous):

and the derivative of e^(-u) is e^(-u) right?

OpenStudy (anonymous):

So it would be 8u^7*e^(-u)+e^(-u)*u^8

OpenStudy (anonymous):

I used to love math

OpenStudy (anonymous):

The derivative of e^(something) is (derivative of something)*e^(something), so \[\frac{d}{dx}(e^{-u}) = -e^{-u}\]Note the negative sign.

OpenStudy (anonymous):

Because of the -1

OpenStudy (anonymous):

right, so your previous answer was almost right, but it was missing that negative sign.

OpenStudy (anonymous):

Awesome

OpenStudy (anonymous):

I think I can do it from here. I really appreciate your help

OpenStudy (anonymous):

No problem!